OpenStudy (anonymous):

What is the 9th term of the geometric sequence 3,12,48,...? Answers to choose from. a)26,244 b)78,732 c)196,608 d)786,432

OpenStudy (anonymous):

@Spectrum

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):

First term =3 Common ratio =12/3 =4

OpenStudy (anonymous):

Now, FIND t9

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

okay not really

OpenStudy (anonymous):

i dont know how to set it up @sauravshakya

OpenStudy (anonymous):

t9 = ar^(9-1) =3*4^(9-1)

OpenStudy (anonymous):

Now, can u?

OpenStudy (anonymous):

i got 26,244

OpenStudy (anonymous):

is this correct @sauravshakya

OpenStudy (anonymous):

Nope, it is c)196,608

OpenStudy (anonymous):

oh i see what i did thank you.

OpenStudy (anonymous):

welcome